OpenStudy (anonymous):

Calculus help!! Each side of a square is increasing at a rate of 6cm/s. At what rate is the area of the square increasing when the area of the square is 16cm2

OpenStudy (kira_yamato):

A = x^2 dx / dt = 6 When A = 16, x = 4 dA/dt = 2x dx/dt dA/dt = 2(4)(6) = 48

OpenStudy (anonymous):

do you understand how dA/dt = 2x dx/dt?

OpenStudy (anonymous):

No I was trying to understand

OpenStudy (anonymous):